David Fifield

Life Ambition Statement

To begin, I will discuss how my experiences in the Boy Scouts will help me in the future, and how I intend to return some of what I have been given. The most important thing I have gained from Scouts is leadership. From having held several leadership positions in the troop, and from having served as a leader on numerous scouting activities, I have skills that I would not have otherwise. I believe that a thorough knowledge of leadership skills will serve me well in the future.

My immediate plans for scouts are to help form the Troop’s Venturing crew and pursue the Ranger award. I also want to offer my services to the Troop as an Assistant Scoutmaster. I look forward to turning over the position of Instructor and training someone else to hold that position. I want to be able to participate more in the administration of the Troop.

Looking into the future, there are several paths I might take. As far as a career, my most valuable skill is computer programming. I am proficient in C and know Perl and Z80 assembly language. Also, I have considerable experience with Unix system administration. Having spoken to advisors and people in the business, I know that technical school or special classes will be sufficient to perform most jobs requiring these skills. Whether I decide to attend technical school or apply to a college, my parents have told me that they will support me.

I hope to employ some of what I have learned in scouts to my future life. I believe that what I know now and what I will earn will make my coming years fruitful and enjoyable.